You
In a demonstration of the photoelectric effect, suppose that a minimum energy of 5*10^-19J (3.1 eV) is required to dislodge an electron from a metal surface. What is the wavelength (in nm) of radiation for which the detector registers a response? (Planks constant h=6.63*10^-34 kgm^2/s)
1
Expert's answer
2020-03-06T10:26:59-0500
"E=\frac{hc}{\lambda}"

"5\cdot 10^{-19}=\frac{(6.63\cdot 10^{-34})(3\cdot 10^{8})}{\lambda}"

"\lambda=4\cdot 10^{-7}m=400\ nm"